# Which Way Does Time Flow? A Psychophysics-Grounded Evaluation for Vision–Language Models

### v2

@@ Abstract
Modern vision–language models (VLMs) excel at many multimodal tasks, yet their grasp of temporal information in video remains weak and, crucially, under-evaluated. We probe this gap with a deceptively simple but revealing challenge: judging the arrow of time (AoT)—whether a short clip is played forward or backward. We introduce AoT-PsyPhyBENCH, a psychophysically validated benchmark that tests whether VLMs can infer temporal direction in natural videos using the same stimuli and behavioral baselines established for humans. Our comprehensive evaluation of open-weight and proprietary, reasoning and non-reasoning VLMs reveals that most models perform near chance, and even the best lag far behind human accuracy on physically irreversible processes (e.g., free fall, diffusion/explosion) and causal manual actions (division/addition) that humans recognize almost instantly. These results highlight a fundamental gap in current multimodal systems: while they capture rich visual–semantic correlations, they lack the inductive biases required for temporal continuity and causal understanding. We release the code and data for AoT-PsyPhyBENCH to encourage further progress in the physical and temporal reasoning capabilities of VLMs.
@@ 1 Introduction
Vision–language models (VLMs) have recently achieved remarkable progress in multimodal understanding, including tasks such as video captioning, retrieval, and question answering bai2025qwen2; qwen_qvq_2024; liu-etal-2024-mibench; wang2024muirbench; xiao2021nextqa; wu2024starbench; fu2025videomme. Yet a fundamental question remains: do these models possess the implicit physical and causal assumptions, i.e. the inductive biases, that humans rely on to interpret the physical world? Among the most basic of such biases is the arrow of time (AoT): the implicit assumption that events unfold irreversibly from past to future, constrained by gravity, entropy, and causality. Judging whether a short video plays forward or backward provides a minimal yet powerful diagnostic of such bias: a system that has internalized temporal irreversibility should recognize when physical regularities are violated, whereas a system that relies only on visual correlation will fail despite extensive training.
Figure 1: Overview of the arrow-of-time (AoT) task. Humans identify the AoT for both forward and backward playbacks with high accuracy; however, VLMs perform substantially worse and exhibit a label-prediction bias, preferring Forward (see Section 6.2).
@@
To anchor this diagnostic in human performance, hanyu2023ready conducted a psychophysical study using 360 three-second natural video clips from everyday activities. They found a global forward bias: participants more often judged reversed clips as forward (39% errors on reversed vs. 9% on forward trials). However, humans detected reversals rapidly and almost flawlessly in five specific categories: Fall (free fall) and Diffusion (diffusion/explosion), which involve physically irreversible processes governed by gravity and entropy; Proceed (forward locomotion); and Division (manual division) and Put (manual addition), which involve agent-driven causal sequences. In these cases, reversed playback becomes visibly implausible. For example, when a big chunk of snow ascends back onto the roof, defying gravity (Figure 1). These results demonstrate that humans exploit not only low-level motion cues but prior knowledge of physical regularities and causal structure, providing a psychophysically validated foundation for testing AoT in VLMs.
@@
Building on this, we introduce AoT-PsyPhyBENCH (pronounced “AoT-sci-fi bench”), an evaluation framework that directly inherits validated stimuli and behavioral baselines from hanyu2023ready. We systematically test modern VLMs, from open-weight to proprietary reasoning/non-reasoning models, on the same AoT judgment task. Our findings reveal that most models perform near chance even in "obvious" irreversible processes, with the best setting lagging humans by approximately 29 percentage points, suggesting that they lack the inductive biases underpinning human causal and temporal perception (Figure 1).
@@
Our contributions are threefold. (1) Comprehensive evaluation. We present the first systematic evaluation of modern VLMs on psychophysically grounded AoT judgment across zero-shot, few-shot, chain-of-thought prompting, and supervised fine-tuning. Despite apparent task simplicity, the gains from elaborate prompting and fine-tuning are minimal and inconsistent, indicating that the bottleneck is temporal and physical understanding, not instruction following. (2) Psychophysically validated benchmark. We introduce AoT-PsyPhyBENCH, curated from hanyu2023ready’s psychophysical dataset by retaining only clips with clear human consensus on irreversible processes while excluding ambiguous cyclic motions. This yields a low-ambiguity benchmark enabling direct, reliable human–model comparison on temporal direction judgment. (3) Open resources. We release our benchmark, evaluation scripts, and model outputs to advance research on temporal reasoning and physical understanding in VLMs. [^1] https://github.com/mattashiho233/AoT-PsyPhyBENCH
@@ 2.1 Vision–Language Models (VLMs)
Modern VLMs couple a vision encoder with a large language model (LLM) via a multimodal connector and are typically aligned through instruction tuning on image–/video–text data wang2024qwen2vl; qwen_qvq_2024; azzolini2025cosmos; bai2025qwen2. Despite rapid progress, frontier VLMs often fail on video tasks that depend on event order xue2025seeing.
@@
Building on this taxonomy, we evaluate the following models. Proprietary non-reasoning: GPT-4o and GPT-4.1 openai_gpt4o_blog_2024; openai_gpt41_blog_2025. Proprietary reasoning: o3, o4-mini, and GPT-5 openai_o3_blog_2025; openai_gpt5_blog_2025; Gemini-2.5-Pro comanici2025gemini. Open-weight non-reasoning: Qwen2-VL wang2024qwen2vl and Qwen2.5-VL bai2025qwen2. Open-weight reasoning: Cosmos-Reason1 azzolini2025cosmos, which combines explicit AoT supervision and reinforcement learning on $\sim$30k forward/reverse video clip pairs with reasoning traces; and QVQ-72B-Preview qwen_qvq_2024.
@@ 2.2 Temporal Reasoning Benchmarks
(1) Lack of temporal dependency. Several recent benchmarks claim to assess temporal reasoning in VLMs, but many do not demonstrate dependence on event order liu-etal-2024-mibench; wang2024muirbench; fu2025videomme; xiao2021nextqa; wu2024starbench. xue2025seeing evaluated strong VLMs on standard temporal benchmarks and found a critical failure: models show minimal or no performance degradation when frames are shuffled or reversed, indicating these tasks can be solved through scene context without temporal understanding. In contrast, the AoT judgment is inherently temporal: Reversed videos violate physical laws, making temporal order task critical, rather than incidental.
(2) Lack of psychophysical validation. Existing AoT benchmarks simply reverse video datasets bagad2023test; wang2023paxion; du2024reversed; agarwal2025cosmos; xue2025seeing without controlled human baselines, for instance, with quantitative measurements of accuracy, reaction times, and category-specific effects. This creates systematic evaluation problems: difficulty remains uncalibrated (videos ambiguous to humans become spurious failure cases), model–human performance gaps cannot be quantified, systematic biases (forward bias, category effects) go undetected, and bidirectionally plausible cases contaminate test sets. Psychophysical grounding is essential to distinguish genuine deficiencies from artifacts of ambiguous stimuli.
@@ 3 AoT-PsyPhyBENCH: A Psychophysically Validated Arrow-of-Time Benchmark
We evaluate whether vision–language models (VLMs) can infer the arrow of time—that is, distinguish forward and backward playbacks in everyday videos. Building on the psychophysical study of hanyu2023ready, enabling direct comparison with human performance. The source dataset comprises 360 three-second clips (29.97 FPS) from the Moments in Time dataset monfort2019moments, covering a broad range of daily dynamics. Ten participants viewed every clip twice—once forward and once reversed—across two sessions on different days (each session: 180 forward, 180 reversed; order randomized per participant). Three raters annotated six motion categories: (1) Proceed: forward locomotion of people, animals, or vehicles; (2) Fall: free-fall/ballistic motion under gravity; (3) Diffusion: centrifugal diffusion or small-particle explosions; (4) Division: division of material by hand or tool; (5) Put: addition/construction of material by hand; and (6) Reciprocal: reciprocating (cyclic) motion (Table 1). The first five categories reflect irreversible processes that, when reversed, tend to violate fundamental physical regularities (entropy increase, gravity, causal ordering), whereas reciprocating motion is bidirectional and often appears plausible in both directions. Categories are not mutually exclusive. Human performance exhibited a pronounced forward bias: participants made errors on 39% of reversed trials versus 9% of forward trials, with accuracy strongly modulated by motions from the first five categories. Performance was significantly lower for the reciprocating motion category.
We introduce AoT-PsyPhyBENCH, a curated subset of hanyu2023ready that excludes the reciprocal motion category (category 6) and retains only high-consensus clips from categories 1–5. This exclusion is justified by human performance: the reciprocal motion category achieved only 61.1% overall accuracy (accuracy with forward and backward videos combined) compared to 83.8–93.5% for the other motion categories, substantially degrading benchmark reliability (77.6% with cyclic motion vs. 89.2% without). From the original 360 clips (720 videos with reversals), our filtering retains 212 clips, resulting in a total of 424 videos. This produces a low-ambiguity, psychophysically validated AoT evaluation benchmark that enables direct and reliable human–model comparison. Table 1 provides an overview of the motion categories with comparative statistics illustrating how AoT-PsyPhyBENCH  filters the original dataset.
@@
Table 1: Motion categories following hanyu2023ready, with human performance, counts, and visual examples. Human F1 scores are reported as forward (F)/backward (B) for each motion category. AoT-PsyPhyBENCH  includes categories 1–5 (all clips) and excludes category 6. marks categories where reversal is typically easy for humans (irreversible processes); marks cyclic, bidirectional motions where reversal is challenging.
@@ 4.3 Reasoning Effort Ablation Settings
Simple CoT. The model is instructed to focus on moving elements and identify visual cues indicating playback direction. We manually curated reasoning traces for each demonstration. Due to space limitations, we omitted the example in this version.
Multi-step CoT. We designed structured instructions with multi-step reasoning traces divided into three stages: Observation (objectively describe visible events without assumptions), Assumption (infer plausible physical or causal sequences), and Conclusion (determine whether observations align with or contradict assumptions to judge playback direction). Due to space limitations, we omitted the prompts in this version. A model output example is shown in Figure 2.
@@ 4.4 Supervised Fine-Tuning (SFT) Settings
We constructed in-domain fine-tuning datasets to test whether explicit temporal direction supervision improves model performance. Note that all videos in hanyu2023ready, AoT-PsyPhyBENCH, and these fine-tuning datasets are drawn from Moments in Time monfort2019moments. We randomly sampled 500, 1,500, and 2,500 clips distributed across its 339 action classes in the original dataset. These clips are entirely distinct from the videos in AoT-PsychBench. Each video was used to create two examples: the original (forward, labeled F) and its temporal reverse (backward, labeled B), yielding 1,000, 3,000, and 5,000 training examples.
@@ 4.6 Hyperparameter Settings
SFT Parameters. We froze the vision encoder and applied LoRA hu2021lora (rank 8) to both the language model and the vision-language bridging MLP layer. We used a maximum sequence length of 2,048 tokens, a batch size of 2 per device with gradient accumulation over 2 steps (effective batch size of 4), a learning rate of $1\times 10^{-5}$, and trained for 5 epochs with a cosine learning rate scheduler and 10% warmup.
@@ 5.2 Few-shot Performance
We test whether few-shot examples elicit in-context learning for AoT (Table 3). Overall, few-shot prompting does not consistently improve performance and often amplifies label prediction bias. GPT-5 shows modest, stable gains (56.1→57.4% accuracy; Backward F1: 26.8→30.9 with 4 shots). In contrast, GPT-4.1 degrades with few-shot prompts (60.0→56.0/55.5% accuracy; Backward F1: 57.4→51.7/49.9). GPT-4o shows severe degradation: accuracy drops (52.6→50.7%) while Backward F1 collapses catastrophically (24.9→4.6/5.1), indicating near-complete failure to detect reversed videos. The open model Qwen2.5-VL-72B exhibits minimal accuracy change (49.5→51.9%) but substantially worsened bias (Backward F1: 38.0→18.7/21.1). These results demonstrate that few-shot prompting provides limited benefits for temporal reasoning: only one model (GPT-5) shows modest improvement, while others either decline or amplify their existing forward bias, suggesting that in-context learning is insufficient to elicit robust AoT understanding.
Figure 2: Left: a backward video clip (category: Put). Top-Right: Qwen2.5VL-72B Multi-step CoT reasoning. Bottom-Right: Gemini-2.5-Pro’s self-explained reasoning trace. Qwen2.5VL-72B correctly identified the event in the scene and made a valid assumption, but failed to observe that the event was reversed which led to an incorrect conclusion. In contrast, Gemini-2.5-Pro correctly detected the reversal of the event in Step 2 based on a valid assumption it made in Step 1.
Setting Shots F. F1 B. F1 Acc.
Qwen2.5VL-72B-instruct
No CoT 2 64.2 18.7 50.2
Simple 2 67.3 3.76 51.2
Multi-step 2 66.7 5.48 50.7
5.3 Reasoning Effort Ablation
@@
5.4 SFT Performance
We fine-tuned Qwen2-7B and Qwen2.5-VL-7B on 1,000, 3,000, and 5,000 training examples to test whether explicit supervision improves AoT performance (Table 6). Fine-tuning failed to improve accuracy, which remained near chance level (~50%) across all training set sizes. While fine-tuning reduced the extreme forward bias present in vanilla models, leading to a more balanced forward/backward F1 scores, it did not yield genuine task competence. Performance plateaued after 1,000 examples, with no gains from additional data. These results suggest that AoT task might require different training approaches beyond conventional SFT.

### v3

@@ Abstract
Modern vision–language models (VLMs) excel at many multimodal tasks, yet their grasp of temporal information in video remains weak and has not been adequately evaluated. We probe this gap with a deceptively simple but revealing challenge: judging the arrow of time (AoT)—whether a short clip is played forward or backward. We introduce AoT-PsyPhyBENCH, a psychophysically validated benchmark that tests whether VLMs can infer temporal direction in natural videos using the same stimuli and behavioral baselines established for humans. Our comprehensive evaluation of open-weight and proprietary, reasoning and non-reasoning VLMs reveals that most models perform near chance, and even the best model lags far behind human accuracy on physically irreversible processes (e.g., free fall, diffusion/explosion) and causal manual actions (division/addition) that humans recognize almost instantly. These results highlight a fundamental gap in current multimodal systems: while they capture rich visual–semantic correlations, they lack the inductive biases required for temporal continuity and causal understanding. We release the code and data for AoT-PsyPhyBENCH to encourage further progress in the physical and temporal reasoning capabilities of VLMs.
@@ 1 Introduction
Vision–language models (VLMs) have recently achieved remarkable progress in multimodal understanding, including tasks such as video captioning, retrieval, and question answering Bai et al. (2025); Qwen Team (2024); Liu et al. (2024); Wang et al. (2024a); Xiao et al. (2021); Wu et al. (2024); Fu et al. (2025). Yet a fundamental question remains: do these models possess the implicit physical and causal assumptions, i.e. the inductive biases, that humans rely on to interpret the physical world? Among the most basic of such biases is the arrow of time (AoT): the implicit assumption that events unfold irreversibly from past to future, constrained by gravity, entropy, and causality. Judging whether a short video plays forward or backward provides a minimal yet powerful diagnostic of such bias: a system that has internalized temporal irreversibility should recognize when physical regularities are violated, whereas a system that relies only on visual correlation will fail despite extensive training.
Figure 1: Overview of the arrow-of-time (AoT) task. Given a video played either forward or backward, the model must predict its temporal direction. Humans identify temporal direction accurately, whereas VLMs perform worse and exhibit a bias toward Forward (Section 6.2). Human Forward/Backward F1 is computed on AoT-PsyPhyBENCH, excluding the Reciprocal (cyclic-motion) category.
@@
To anchor this diagnostic in human performance, Hanyu et al. (2023) conducted a psychophysical study using 360 three-second natural video clips from everyday activities. They found a global forward bias: participants more often judged reversed clips as forward (39% errors on reversed vs. 9% on forward trials). However, humans detected reversals rapidly and almost flawlessly in five specific categories: Fall (free fall) and Diffusion (diffusion/explosion), which involve physically irreversible processes governed by gravity and entropy; Proceed (forward locomotion); and Division (manual division) and Put (manual addition), which involve agent-driven causal sequences. In these cases, reversed playback becomes visibly implausible. For example, when a big chunk of snow ascends back onto the roof, defying gravity (Figure 1). These results demonstrate that humans exploit not only low-level motion cues but prior knowledge of physical regularities and causal structure, providing a psychophysically validated foundation for testing AoT in VLMs.
@@
Building on this, we introduce AoT-PsyPhyBENCH (pronounced “AoT-sci-fi bench”), an evaluation framework that directly inherits validated stimuli and behavioral baselines from Hanyu et al. (2023). We systematically test modern VLMs, from open-weight to proprietary reasoning/non-reasoning models, on the same AoT judgment task. Our findings reveal that most models perform near chance even in "obvious" irreversible processes, with the best setting lagging humans by approximately 29 percentage points, suggesting that they lack the inductive biases underpinning human causal and temporal perception (Figure 1).
@@
Our contributions are threefold. (1) Comprehensive evaluation. We present the first systematic evaluation of modern VLMs on psychophysically grounded AoT judgment across zero-shot, few-shot, chain-of-thought prompting, and supervised fine-tuning. Despite apparent task simplicity, the gains from elaborate prompting and fine-tuning are minimal and inconsistent, indicating that the bottleneck is temporal and physical understanding, not instruction following. (2) Psychophysically validated benchmark. We introduce AoT-PsyPhyBENCH, curated from Hanyu et al.’s psychophysical dataset by retaining only clips with clear human consensus on irreversible processes while excluding ambiguous cyclic motions. This yields a low-ambiguity benchmark enabling direct, reliable human–model comparison on temporal direction judgment. (3) Open resources. We release our benchmark, evaluation scripts, and model outputs to advance research on temporal reasoning and physical understanding in VLMs. [^1] https://mattashiho233.github.io/AoT-PsyPhyBENCH-website/
@@ 2.1 Vision–Language Models (VLMs)
Modern VLMs couple a vision encoder with a large language model (LLM) via a multimodal connector and are typically aligned through instruction tuning on image–/video–text data Wang et al. (2024b); Qwen Team (2024); Azzolini et al. (2025); Bai et al. (2025). Despite rapid progress, frontier VLMs often fail on video tasks that depend on event order Xue et al. (2025).
@@
Building on this taxonomy, we evaluate the following models. Proprietary non-reasoning: GPT-4o and GPT-4.1 OpenAI (2024, 2025a). Proprietary reasoning: o3, o4-mini, and GPT-5 OpenAI (2025c, b); Gemini-2.5-Pro Comanici et al. (2025). Open-weight non-reasoning: Qwen2-VL Wang et al. (2024b) and Qwen2.5-VL Bai et al. (2025). Open-weight reasoning: Cosmos-Reason1 Azzolini et al. (2025), which combines explicit AoT supervision and reinforcement learning on $\sim$30k pairs of forward/reverse video clips with reasoning traces; and QVQ-72B-Preview Qwen Team (2024).
@@ 2.2 Temporal Reasoning Benchmarks
(1) Lack of temporal dependency. Several recent benchmarks claim to assess temporal reasoning in VLMs, but many do not demonstrate dependence on event order Liu et al. (2024); Wang et al. (2024a); Fu et al. (2025); Xiao et al. (2021); Wu et al. (2024). Xue et al. (2025) evaluated strong VLMs on standard temporal benchmarks and found a critical failure: models show minimal or no performance degradation when frames are shuffled or reversed, indicating these tasks can be solved through scene context without temporal understanding. In contrast, the AoT judgment is inherently temporal: reversed videos violate physical laws, making temporal order task critical, rather than incidental.
(2) Lack of psychophysical validation. Existing AoT benchmarks simply reverse video datasets Bagad et al. (2023); Wang et al. (2023); Du et al. (2024); Agarwal et al. (2025); Xue et al. (2025) without controlled human baselines, for instance, with quantitative measurements of accuracy, reaction times, and category-specific effects. This creates systematic evaluation problems: difficulty remains uncalibrated (videos ambiguous to humans become spurious failure cases), model–human performance gaps cannot be quantified, systematic biases (forward bias, category effects) go undetected, and bidirectionally plausible cases contaminate test sets. Psychophysical grounding is essential to distinguish genuine deficiencies from artifacts of ambiguous stimuli.
@@ 3 AoT-PsyPhyBENCH: A Psychophysically Validated Arrow-of-Time Benchmark
We evaluate whether vision–language models (VLMs) can infer the arrow of time, that is, distinguish between forward and backward playbacks in everyday videos. Our benchmark builds on the psychophysical study of Hanyu et al. (2023), enabling direct comparison with human performance. The source dataset comprises 360 three-second clips (29.97 FPS) from the Moments in Time dataset Monfort et al. (2019), covering a broad range of daily dynamics. Ten participants viewed every clip twice—once forward and once reversed—across two sessions on different days (each session: 180 forward, 180 reversed; order randomized per participant). Three raters annotated six motion categories: (1) Proceed: forward locomotion of people, animals, or vehicles; (2) Fall: free-fall/ballistic motion under gravity; (3) Diffusion: centrifugal diffusion or small-particle explosions; (4) Division: division of material by hand or tool; (5) Put: addition/construction of material by hand; and (6) Reciprocal: reciprocating (cyclic) motion (Table 1). The first five categories reflect irreversible processes that, when reversed, tend to violate fundamental physical regularities (entropy increase, gravity, causal ordering), whereas reciprocating motion is bidirectional and often appears plausible in both directions. Categories are not mutually exclusive. Human performance exhibited a pronounced forward bias: participants made errors on 39% of reversed trials versus 9% of forward trials, with accuracy strongly modulated by motions from the first five categories. Performance was significantly lower for the reciprocating motion category.
We introduce AoT-PsyPhyBENCH, a curated subset of Hanyu et al. (2023) that excludes the reciprocal motion category (category 6) and retains only high-consensus clips from categories 1–5. This exclusion is justified by human performance: the reciprocal motion category achieved only 61.1% overall accuracy (accuracy with forward and backward videos combined) compared to 81.3–85.1% for the other motion categories, substantially degrading benchmark reliability (77.6% with cyclic motion vs. 89.0% without). From the original 360 clips (720 videos with reversals), our filtering retains 212 clips, resulting in a total of 424 videos. This yields a low-ambiguity benchmark that enables direct and reliable comparison between humans and models. Table 1 provides an overview of the motion categories with comparative statistics illustrating how AoT-PsyPhyBENCH  filters the original dataset.
@@
Table 1: Motion categories following Hanyu et al. (2023), with human performance, counts, and visual examples. Human F1 scores are reported as forward (F)/backward (B) for each motion category. AoT-PsyPhyBENCH  includes categories 1–5 (all clips) and excludes category 6. marks categories where reversal is typically easy for humans (irreversible processes); marks cyclic, bidirectional motions where reversal is challenging.
@@ 4.3 Reasoning Effort Ablation Settings
Figure 2: Manually curated chain-of-thought reasoning traces for a Forward example.
Simple CoT. The model is instructed to focus on moving elements and identify visual cues indicating playback direction. We manually curated reasoning traces for each demonstration. The instruction for this setting is provided in Figure 6. Two manually curated reasoning traces are shown in Figure 23.
Multi-step CoT. We designed structured instructions with multi-step reasoning traces divided into three stages: Observation (objectively describe visible events without assumptions), Assumption (infer plausible physical or causal sequences), and Conclusion (determine whether observations align with or contradict assumptions to judge playback direction). The instruction for this setting is provided in Figure 7. Two manually curated reasoning traces are shown in Figure 23. A model output example is shown in Figure 4.
Figure 3: Manually curated chain-of-thought reasoning traces for a Backward example.
@@ 4.4 Supervised Fine-Tuning (SFT) Settings
We constructed in-domain fine-tuning datasets to test whether explicit temporal direction supervision improves model performance. Note that all videos in Hanyu et al. (2023), AoT-PsyPhyBENCH, and these fine-tuning datasets are drawn from Moments in Time Monfort et al. (2019). We randomly sampled 500, 1,500, and 2,500 clips distributed across its 339 action classes in the original dataset. These clips are entirely distinct from the videos in AoT-PsychBench. Each video was used to create two examples: the original (forward, labeled F) and its temporal reverse (backward, labeled B), yielding 1,000, 3,000, and 5,000 training examples.
@@ 4.6 Hyperparameter Settings
SFT Parameters. We froze the vision encoder and applied LoRA Hu et al. (2021) (rank 8) to both the language model and the vision-language bridging MLP layer. We used a maximum sequence length of 2,048 tokens, a batch size of 2 per device with gradient accumulation over 2 steps (effective batch size of 4), a learning rate of $1\times 10^{-5}$, and trained for 5 epochs with a cosine learning rate scheduler and 10% warmup.
@@ 5 Results
Model #Shots F. F1 B. F1 Acc.
Open Model
Qwen2.5 VL-72B 0 57.4 38.0 49.5
2 64.2 18.7 50.2
4 65.4 21.1 51.9
Proprietary Models
GPT-4.1 0 62.3 57.4 60.0
2 59.5 51.7 56.0
4 60.0 49.9 55.5
GPT-4o 0 65.4 24.9 52.6
2 66.8 4.61 50.7
4 66.5 5.07 50.7
GPT-5 0 68.7 26.8 56.1
2 69.5 29.2 57.4
4 69.2 30.9 57.4
Table 3: Few-shot performance.
Model Effort F. F1 B. F1 Acc.
Proprietary Models
GPT-5 Low 68.9 34.4 57.8
Medium 68.7 26.8 56.1
High 69.1 25.7 56.4
Gemini- 2.5-pro Low 67.2 52.1 61.1
Medium 65.9 51.4 59.9
High 65.3 49.0 58.7
Table 4: Reasoning effort ablation.
Setting Shots F. F1 B. F1 Acc.
Qwen2.5VL-72B-instruct
No CoT 2 64.2 18.7 50.2
Simple 2 67.3 3.76 51.2
Multi-step 2 66.7 5.48 50.7
Data Size F. F1 B. F1 Acc.
Qwen2-7B Training
Vanilla 66.7 0.0 50.0
1000 64.1 12.2 49.1
3000 54.2 48.2 51.4
5000 59.4 31.6 49.1
Qwen2.5VL-7B Training
Vanilla 63.0 19.5 49.3
1000 54.8 43.5 49.8
3000 54.5 43.4 49.5
5000 54.5 43.4 49.5
Table 6: SFT performance.
Model FPS F. F1 B. F1 Acc.
Human 30 90.0 88.0 89.2
GPT-4.1 2 62.1 51.2 57.3
4 62.5 57.4 60.1
8 62.9 55.8 59.7
16 62.4 53.7 58.5
30 64.4 55.9 60.6
Table 7: FPS ablation results.
@@ 5.2 Few-shot Performance
We test whether few-shot examples elicit in-context learning for AoT (Table 8). Overall, few-shot prompting does not consistently improve performance and often amplifies label prediction bias. GPT-5 shows modest, stable gains (Accuracy: 56.1→57.4/57.4%; Forward F1: 68.7→69.5/69.2; Backward F1: 26.8→29.2/30.9). In contrast, GPT-4.1 degrades with few-shot prompts (Accuracy: 60.0→56.0/55.5%; Forward F1: 62.3→59.5/60.0; Backward F1: 57.4→51.7/49.9). GPT-4o shows severe degradation: accuracy drops (52.6→50.7%), with Backward F1 collapsing catastrophically (24.9→4.6/5.1), indicating near-complete failure to detect reversed videos. The open model Qwen2.5-VL-72B exhibits minimal accuracy change (49.5→51.9%) but substantially worsened bias (Backward F1: 38.0→18.7/21.1). These results demonstrate that few-shot prompting provides limited benefits for temporal reasoning: only one model (GPT-5) shows modest improvement, while others either decline or amplify their existing forward bias, suggesting that in-context learning is insufficient to elicit robust AoT understanding.
@@
Setting Shots F. F1 B. F1 Acc.
Qwen2.5VL-72B-instruct
No CoT 2 64.2 18.7 50.2
Simple 2 67.3 3.76 51.2
Multi-step 2 66.7 5.48 50.7
Figure 4: Left: a backward video clip (category: Put). Top-Right: Qwen2.5VL-72B Multi-step CoT reasoning. Bottom-Right: Gemini-2.5-Pro’s self-explained reasoning trace. Qwen2.5VL-72B correctly identified the event in the scene and made a valid assumption, but failed to observe that the event was reversed which led to an incorrect conclusion. In contrast, Gemini-2.5-Pro correctly detected the reversal of the event in Step 2 based on a valid assumption it made in Step 1. The Gemini-2.5-Pro output is from the low reasoning effort setting.
5.3 Reasoning Effort Ablation